OpenStudy (anonymous):

you randomly select 3 different cards from a deck. find the probability of selecting 3 hearts

OpenStudy (kropot72):

\[P(3\ hearts)=\frac{13C3}{52C3}=\frac{13!}{3!10!}\times \frac{3!49!}{52!}=you\ can\ calculate\]